OpenStudy (anonymous):

Simplify the expression 12x^-4y^2/3x^2y^-4

OpenStudy (jdoe0001):

\(\Large \bf { a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}} \\ \quad \\ \quad \\ \cfrac{12x^{-4}y^2}{3x^2y^{-4}}\implies \cfrac{12\cdot \frac{1}{x^{\color{red}{ 4}}}\cdot y^2}{3x^2\cdot \frac{1}{y^{\color{red}{ 4}}}}\implies \cfrac{\frac{12y^2}{x^4}}{\frac{3x^2}{y^4}} \\ \quad \\ recall\implies {\color{brown}{ \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}}} }\)